A 1840 kg car has just passed a diner. The closest the car came to the diner was R = 24.5 m, but the car has driven another x = 38.7 m in the 2.45 s elapsed since then. If the car travels at constant speed, what is the current magnitude of the angular momentum of the car about the diner?
Answer #1

here,

mass of the car , m = 1840 kg

x = 38.7 m , t = 2.45 s

speed of car , v= x/t

v = 15.8 m/s

R = 24.5 m

the current magnitude of the angular momentum of the car about the diner , L = m * v * r

L = 1840 * 15.8 * 24.5

L = 7.12 * 10^5 kg.m^2/s

the current magnitude of the angular momentum of the car about the diner is 7.12 * 10^5 kg.m^2/s